calculate the air: O2 ratio of 30% O2: air entrainment device ___________
Answer #1

Calculating the above using the below formula:

- Air/O2 ratio= 100- O2%/ O2%- 21

= 100-30/ 30-21

= 70/ 9

= 7.7 = 8 = 8:1

Air : oxygen ratio is 8:1